1. (10 points) Find the parametric equation for a Bezier curve with five control points Po, P1, P2, P3, and P4.
Answer #1

HT Solution P.N. I Equation for a Bezier Cusre Standard equation at Bezier Cuę ve 4-Con pol point Plo) = !0°(1-05 po t (? )u(P.N.12. Standard egn is (4-Control points) P(4)= (1-01? Po + 34(1-0)2p, +342(1-4)P+U3P3 + For 2 Coordinate P2 (0) = (1-0) Poz3 The Can be Calculated some was But hege yzo in in all point hence P(u)y=0 U= 0, 0.2, 0.6 0.8, 100 U = 0 0.2 PY 0.6 0.8 30 1